# Breaking the Dimensional Barrier: A Pontryagin-Guided Direct Policy Optimization for Continuous-Time Multi-Asset Portfolio

### v2

@@ 3.3.1 The Parabolic Nature of Financial Models and PDE Regularity
Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Two-Stage PG-DPO method. The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where the Two-Stage PG-DPO effectively leverages the expected regularity of the Costates. The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
3.3.2 Mathematical Foundation: Derivative Stability via PDE Comparison
A key theoretical insight, particularly demonstrable in well-structured (e.g., affine) models, comes from analyzing the stability of value function derivatives under policy perturbations using regularity theory for parabolic partial differential equations (e.g., Evans, , 2022; Krylov, , 2024). This perspective illustrates how closeness of policies (in a specific sense) can imply closeness of derivatives, providing a foundation for why the Two-Stage approach might be effective.
First, let us recall the PDE associated with a given policy. For an admissible and sufficiently smooth policy $(\pi(t,z),c(t,z))$, where $z=(x,y)$ represents the state, the corresponding value function $V(t,z)$ (representing the expected utility achieved under this *specific* policy from state $z$ at time $t$) satisfies the linear parabolic PDE:
$-\frac{\partial V}{\partial t}=\mathcal{L}_{\pi,c}[V](t,z)+f_{\pi,c}(t,z),\quad\text{subject to }V(T,z)=g(z).$ (16)
Here, $g(z)=\kappa e^{-\delta T}U(x)$ is the terminal utility, $f_{\pi,c}(t,z)=e^{-\delta t}U(c(t,z))$ is the instantaneous utility source term under policy $c$, and $\mathcal{L}_{\pi,c}$ is the second-order linear differential generator associated with the state dynamics under policy $(\pi,c)$:
$\mathcal{L}_{\pi,c}[V]=$ $V_{x}\left[x\left((1-\pi^{\top}\mathbf{1})r(t,y)+\pi^{\top}\mu(t,y)\right)-c(t,z)\right]+V_{\mathbf{y}}^{\top}\mu_{Y}(t,y)$
$+\frac{1}{2}V_{xx}x^{2}\pi^{\top}\sigma(t,y)\Psi\sigma(t,y)^{\top}\pi$
$+V_{x\mathbf{y}}^{\top}x\pi^{\top}\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$ (17)
$+\frac{1}{2}\text{Tr}\left(V_{\mathbf{y}\mathbf{y}}\sigma_{Y}(t,y)\Phi\sigma_{Y}(t,y)^{\top}\right)-\delta V$
(We incorporate the time discount term $-\delta V$ into the operator $\mathcal{L}_{\pi,c}$ for convenience). Let $V^{*}(t,z)$ denote the true optimal value function of the control problem defined in Section 2.1, achieved by the optimal policy $(\pi^{*}(t,z),c^{*}(t,z))$. Assuming sufficient regularity, $V^{*}$ satisfies the non-linear Hamilton-Jacobi–Bellman (HJB) equation (Fleming & Soner, , 2006), which implies it also satisfies the linear PDE (16) when the optimal policy $(\pi^{*},c^{*})$ is used:
$-\frac{\partial V^{*}}{\partial t}=\mathcal{L}_{\pi^{*},c^{*}}[V^{*}](t,z)+f_{\pi^{*},c^{*}}(t,z),\quad\text{subject to }V^{*}(T,z)=g(z).$ (18)
To formalize the notion of policy closeness relevant to the PDE coefficients, we introduce the following definition. Let $\Sigma_{\text{asset}}(t,y)=\sigma(t,y)\Psi\sigma(t,y)^{\top}$ and $\Sigma_{\text{asset-state}}(t,y)=\sigma(t,y)\rho\sigma_{Y}(t,y)^{\top}$.
Definition 1 (Coefficient-relevant $\epsilon$-closeness).
Let the spatial domain be
$\Omega\subseteq(0,\infty)\times\mathbb{R}^{k},\qquad z=(x,\mathbf{y})\in\Omega,$
i.e. $x$ is wealth and $\mathbf{y}\in\mathbb{R}^{k}$ is the $k$-dimensional state vector. (In practice $\Omega$ can be either the full space or a bounded truncated box that contains the region the process visits with high probability.)
Fix exponents $p>d+2$ and $q\geq 2$, where $d=\dim\Omega=k+1$. For any measurable function $h(t,z)$ set
[…]
Let $(\pi^{*},c^{*})$ be the optimal policy and $(\pi,c)$ any admissible policy. We say $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ on $Q_{T}=(0,T)\times\Omega$ if
$\|c-c^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\mathbf{1}-\pi^{*\top}\mathbf{1}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\bm{\mu}-\pi^{*\top}\bm{\mu}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\Sigma_{\text{asset}}\pi-\pi^{*\top}\Sigma_{\text{asset}}\pi^{*}\|_{L^{q,p}}$ $\leq\epsilon,$
$\|\pi^{\top}\Sigma_{\text{asset-state}}-\pi^{*\top}\Sigma_{\text{asset-state}}\|_{L^{q,p}}$ $\leq\epsilon.$
Proposition 3.1 (Derivative closeness under coefficient $\epsilon$-closeness).
Let $V^{*}$ be the optimal value function corresponding to the optimal policy $(\pi^{*},c^{*})$ (cf. (18)) and let $V$ be the value function generated by another admissible policy $(\pi,c)$ (cf. (16)). Set
$D\;:=\;V^{*}-V,\qquad Q_{T}:=(0,T)\times\Omega.$
Assume
  1. (A1)
    Uniform parabolicity. The operator $\mathcal{L}_{\pi^{*},c^{*}}$ is uniformly parabolic on $Q_{T}$.
  2. (A2)
    Baseline regularity. There exist exponents $p>d+2$ and $q\geq 2$ (with $d=\dim\Omega=k+1$) such that the value functions lie in the Sobolev space (e.g., Adams & Fournier, , 2003; Evans, , 2022)
    $V^{*},\,V\;\in\;L^{q}\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}.$
  3. (A3)
    Coefficient $\epsilon$-closeness. The candidate policy $(\pi,c)$ is $(\epsilon,p,q)$-close to $(\pi^{*},c^{*})$ in the sense of Definition 1.
Then $D$ satisfies the linear PDE
$-\partial_{t}D\;=\;\mathcal{L}_{\pi^{*},c^{*}}[D]\;+\;S^{\prime},\quad\text{where}\quad S^{\prime}=(\mathcal{L}_{\pi^{*},c^{*}}-\mathcal{L}_{\pi,c})[V]+\bigl{(}f_{\pi^{*},c^{*}}-f_{\pi,c}\bigr{)}.$
Moreover, the source term is small in the weighted norm:
$\|S^{\prime}\|_{L^{q,p}}\;\leq\;C_{0}\,\epsilon,$
and by standard $L^{p}$ / Schauder estimates for uniformly parabolic equations (see, e.g., Evans, , 2022, Ch. 7),
$\|D\|_{L^{q}\!\bigl{(}0,T;W^{2,p}(\Omega)\bigr{)}}\;\leq\;C\,\|S^{\prime}\|_{L^{q,p}}\;\leq\;C\,\epsilon,$
for a constant C independent of $\epsilon$.
Consequently, the value-function gap and all first- and second-order spatial derivatives satisfy
$\bigl{\|}V^{*}-V\bigr{\|}_{L^{q,p}}+\bigl{\|}\nabla V^{*}-\nabla V\bigr{\|}_{L^{q,p}}+\bigl{\|}D^{2}V^{*}-D^{2}V\bigr{\|}_{L^{q,p}}\;\leq\;C\,\epsilon.$
Hence the costates $(\nabla V)$ produced under $(\pi,c)$ are $O(\epsilon)$-close to the optimal costates on the entire space–time domain.
Remark on assumptions, PG–DPO rationale, and scope of the result.
Proposition 3.1 makes precise the intuition for the uniformly–parabolic class (“U class”’): if two policies yield close PDE-coefficients in the weak sense of Definition 1 (only a global [^1] A still weaker occupation-weighted norm is discussed below. $L^{q,p}$-distance), then their value functions—and hence their gradients (costates)—differ by at most $O(\epsilon)$. Assumptions (A1)–(A2) are the classical entry ticket for $L^{p}$ / Schauder regularity (Evans, , 2022) and are easily verified for affine models or for models that become uniformly elliptic after adding a small floor to each diffusion coefficient (e.g. replace $\sigma(t,z)$ by $\sqrt{\sigma(t,z)^{2}+\varepsilon^{2}}$ with $\varepsilon\sim 10^{-4}$). Such a floor leaves the economics virtually unchanged while preventing the diffusion matrix from collapsing to zero, which is exactly what uniform parabolicity requires. The genuinely new ingredient is (A3), the $\epsilon$-coefficient closeness that Stage 1 of PG-DPO is designed to enforce.
Stage 1 produces a sub-optimal policy $(\pi,c)$ whose coefficient error is $\epsilon$. Via Proposition 3.1 this translates into an $O(\epsilon)$ error for the costate $\lambda\approx\nabla V$, so the BPTT outputs obtained during the short warm-up are already $O(\epsilon)$-accurate surrogates of the true optimal costates. Stage 2 simply substitutes these estimates into the PMP first-order conditions (Pontryagin, , 2014; Yong & Zhou, , 2012), thereby constructing a near-optimal control without having to train the policy network to full convergence.
Global vs.– local norms.– The result is stated with a global $\|\,\cdot\,\|_{L^{q,p}}$ norm for transparency. If desired, one may replace it by the occupation-weighted norm
[…]
where $\mu_{(\pi,c)}$ is the state-density under the candidate policy. Because the proof relies on local PDE estimates, the same $O(\epsilon)$ bound holds—now only on the subset of the state space that the process actually visits with non-negligible probability. This is valuable for models whose coefficients degenerate in remote regions.
Beyond the U class. – For hypo-elliptic (H class) (cf. Hörmander, , 1967) or boundary-degenerate (D class) models (e.g., Oleinik, , 2012), the full global $L^{p}$ theory is subtler, yet the algorithmic logic of Two-Stage PG-DPO is unaffected:
  • the PMP equations and BPTT apply to any differentiable dynamics;
  • costates typically stabilise faster than the full optimal control;
  • a brief warm-up that makes the coefficient error small where it matters is enough to obtain high-quality costate estimates.
Even when the rigorous guarantee of Proposition 3.1 is unavailable, this costate-first, PMP-second recipe remains an effective practical strategy—an observation confirmed empirically in the numerical section (Section 4) and worth exploring further for genuinely non-affine or degenerate models.
3.3.3 Qualitative Rationale and Broader Applicability
While the formal analysis in Proposition 3.1 provides rigorous support under specific regularity conditions (like uniform parabolicity, often associated with affine models), the core ideas underpinning the Two-Stage PG-DPO approach possess a broader qualitative appeal that suggests applicability beyond these settings.
Crucially, neither the Pontryagin Maximum Principle (PMP) itself (Pontryagin, , 2014; Yong & Zhou, , 2012) nor the mechanics of backpropagation-through-time (BPTT) (e.g., Goodfellow, , 2016) are inherently restricted to affine or uniformly parabolic systems. PMP provides necessary optimality conditions for general nonlinear dynamics, provided the required functions are differentiable. Similarly, BPTT serves as a general algorithm for computing gradient estimates through complex computational graphs derived from simulations, regardless of the underlying model’s linearity.
Furthermore, there is a compelling economic and mathematical intuition that the PMP costates ($\lambda\approx\nabla V$, representing marginal values) might exhibit properties that make them more amenable to estimation than the potentially complex optimal policy $\pi^{*}$ itself, especially in high-dimensional problems. For instance, the costate functions might be representable by smoother or lower-complexity functions compared to $\pi^{*}$. Alternatively, the process of learning or estimating these costate-related functions via BPTT with neural networks might converge more rapidly or stably than directly learning the intricate mapping defined by the optimal policy.
The Two-Stage PG-DPO strategy is designed precisely to leverage this potential advantage. It effectively decouples the optimization challenge into two parts: (i) tackling the potentially more tractable task of estimating costates (and their relevant derivatives) using BPTT during a warm-up phase (ii) employing the robust analytical structure of the PMP first-order conditions to translate these estimates into effective control actions during deployment.
Therefore, even in scenarios where the rigorous derivative closeness guarantees of Proposition 3.1 are harder to establish formally (e.g., for non-affine models, or systems falling into the hypo-elliptic or degenerate parabolic classes), this underlying qualitative rationale remains persuasive. The potential for BPTT-derived costate estimates to stabilize relatively quickly and capture essential optimality information offers a strong motivation for exploring the Two-Stage PG-DPO approach in these more complex settings.
This rationale, significantly reinforced by the numerical success demonstrated in the affine benchmark case (Section 4), positions PG-DPO (particularly the superior Two-Stage variant) as a promising and potentially powerful tool for tackling challenging high-dimensional, possibly non-affine, continuous-time control problems prevalent in finance and other domains. Naturally, its practical effectiveness in such extended settings must be carefully evaluated empirically on a case-by-case basis, highlighting an important direction for future research.
@@ 4 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Two-Stage variants). We focus on the multi-asset, multi-factor Ornstein-Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2. The objective is to maximize the expected utility of terminal wealth, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function with risk aversion $\gamma=2.0$. This specific affine model setting allows for comparison against a semi-analytical benchmark solution derived from the HJB equation (details in Appendix B.2). We compare the policies generated by Baseline PG-DPO and Two-Stage PG-DPO against this benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors. For complete details on the specific parameter values, PG-DPO implementation (network architecture, training setup), benchmark computation, and evaluation procedures, please refer to Appendix C. The subsequent subsections present and analyze the results.
@@ 4.1 Convergence Speed and Accuracy Analysis
This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO methodologies across various dimensional settings. Figure 1 displays the average Root Mean Squared Error (RMSE), averaged across $k=10$ evaluation slices, between the learned policies (either $\bm{\pi}_{\theta}$ for Baseline PG-DPO or $\bm{\pi}^{\mathrm{PMP}}$ for Two-Stage PG-DPO) and the benchmark optimal policy ($\bm{\pi}^{*}$). The RMSE is plotted as a function of training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
@@
Figure 1: Average policy RMSE comparison between Baseline PG-DPO (blue) and Two-Stage PG-DPO (orange) relative to the benchmark solution across different dimensions (k=10 fixed). Values averaged over k evaluation slices. Note the logarithmic scale on the y-axis.
The RMSE for the Two-Stage PG-DPO (orange line in Figure 1) represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ constructed by substituting the BPTT-derived costate estimates into the PMP first-order condition (Eq. (10)). A key advantage is that the Two-Stage approach explicitly combines the analytically distinct myopic and hedging components. As established theoretically and confirmed numerically (see Section 4.2), the error of the myopic component is typically negligible (e.g., on the order of $10^{-7}$). Therefore, the total RMSE for the Two-Stage policy is almost entirely attributable to the error in estimating the hedging component. This highlights the proficiency of the Two-Stage method in accurately capturing hedging demands using BPTT-derived costate derivatives ($\partial_{\mathbf{Y}}\lambda_{t}^{*}$).
Table 2: Summary of Minimum Achieved Policy RMSE and Convergence Epochs
@@
(n assets, k factors) (vs. Benchmark)
n=1, k=10 Baseline PG-DPO $4.27\times 10^{-2}$ $2,600$
2-PG-DPO $2.80\times 10^{-4}$ $1,500$
n=10, k=10 Baseline PG-DPO $1.58\times 10^{-1}$ $10,000$
2-PG-DPO $4.89\times 10^{-3}$ $800$
n=50, k=10 Baseline PG-DPO $1.26\times 10^{-1}$ $100,000$
2-PG-DPO $1.32\times 10^{-2}$ $3,200$
@@
Note: Minimum Policy RMSE values are averaged across the $k=10$ evaluation slices. Epochs indicate the recorded iteration number when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
The key finding from Figure 1 and summarized in Table 2 is the consistent and significant outperformance of the Two-Stage PG-DPO method. Table 2 shows that Two-Stage PG-DPO achieves minimum RMSE values dramatically lower than the Baseline method across all dimensions—approximately 150 times lower for $n=1$, 30 times lower for $n=10$, and nearly 10 times lower for $n=50$. Furthermore, 2-PG-DPO typically reaches its minimum error much earlier in the training process (e.g., at epoch 1,500 vs. 2,600 for $n=1,k=10$, and epoch 800 vs. 10,000 for $n=10,k=10$). Notably, while the convergence time for Baseline PG-DPO appears to increase substantially with the problem dimension (n), the time required for 2-PG-DPO to reach its minimum error seems relatively less sensitive to the number of assets. Even in the high-dimensional $n=50,k=10$ setting, 2-PG-DPO finds a near-optimal solution with an RMSE of $1.32\times 10^{-2}$ around epoch 3200, whereas the Baseline method’s minimum RMSE of $1.26\times 10^{-1}$ is reached only at the end of a much longer training period (epoch 100,000).
This clearly demonstrates the superiority of the Two-Stage approach and strongly supports the rationale discussed in Section 3.3: leveraging the PMP structure via BPTT-estimated costates is more efficient and accurate than end-to-end policy approximation, especially as dimensionality increases. The subsequent section presents a detailed visual analysis for the most challenging high-dimensional case ($n=50,k=10$) to further illustrate the performance difference. Hereafter, we often abbreviate Two-Stage PG-DPO as 2-PG-DPO.

### v3

@@ 3.3 Rationale for the Projected PG-DPO (P-PGDPO) Approach
Models falling into classes (U) and (H) are expected to possess sufficiently regular Costates, providing a favorable setting for the estimation required by the Projected PG-DPO (P-PGDPO) method. The numerical experiments in this paper are conducted using models belonging to classes (U) or (H), where P-PGDPO effectively leverages the expected regularity of the Costates. The method is particularly effective as it generally avoids visiting degenerate boundaries during practical implementation.
Building upon this expected regularity, a more formal mathematical justification for the P-PGDPO approach can be provided. The objective is to establish that if the policy $(\pi,c)$ obtained from the Stage 1 warm-up phase satisfies a sufficiently small Pontryagin first-order condition (FOC) residual $\varepsilon$, then the feedback policy $\widehat{\pi}$ constructed in Stage 2 using BPTT estimates $\widehat{\lambda}$ (with BPTT estimation error $\delta_{\mathrm{BPTT}}$) will be close to the true optimal policy $\pi^{*}$. This analysis relies on certain standing assumptions concerning parabolic regularity and the FOC gap, which are formally stated as Assumption 6 and 7 in Appendix C. We use $L^{q,p}(Q_{T})$ to denote the standard mixed space-time Lebesgue norm over the domain $Q_{T}=(0,T)\times\Omega\subseteq\mathbb{R}^{d+1}$, as detailed in Appendix C.
The main result of this analysis is encapsulated in the following theorem, the proof of which is detailed in Appendix C along with supporting lemmas and propositions.
Theorem 1 (Policy Gap Bound).
Under Assumptions 6 and 7 (detailed in Appendix C), let $(\pi,c)$ be the policy from Stage 1 satisfying the FOC-gap condition $\|\mathcal{R}_{\mathrm{FOC}}(\pi,c;\lambda)\|_{L^{q,p}}\leq\varepsilon$, where $\lambda=\nabla V^{\pi,c}$. Let $\pi^{*}$ be the true optimal policy, and let $\widehat{\pi}=F(\widehat{\lambda})$ be the policy deployed in Stage 2 using BPTT estimates $\widehat{\lambda}$, where $F$ is the explicit mapping from costates $(\lambda,\nabla\lambda)$ to controls derived from the PMP FOCs. Assume $F$ is locally Lipschitz continuous with constant $C_{F}$. If the BPTT estimation error $\|\widehat{\lambda}-\nabla V^{\pi,c}\|_{L^{q,p}}\leq\delta_{\mathrm{BPTT}}$ (where $\delta_{\mathrm{BPTT}}$ quantifies the precision of BPTT-derived costate estimates relative to Stage 1 costates), then the policy gap is bounded by:
$\boxed{\|\widehat{\pi}-\pi^{*}\|_{L^{q,p}}\;\leq\;C_{\mathrm{tot}}(\delta_{\mathrm{BPTT}}+\varepsilon)}$
Here, $C_{\mathrm{tot}}$ is a constant that depends on $C_{F}$ and the constant $C_{v}^{\prime\prime}$ (derived from Proposition C.2 in Appendix C).
The proof of Theorem 1 follows a logical chain: the FOC-gap $(\leq\varepsilon)$ implies that errors in PDE coefficients and source terms are of order $O(\varepsilon)$. These, in turn, through PDE stability, imply that the error in the value function’s gradient $E_{v}=\|\nabla V^{\pi,c}-\nabla V^{*}\|_{L^{q,p}}$ is also of order $O(\varepsilon)$. Combining this with the BPTT estimation error $\delta_{\mathrm{BPTT}}$ leads to the final policy-gap bound. The detailed arguments, including the necessary lemmas and propositions, are provided as sketches in Appendix C. It is important to note that all constants involved depend quantitatively on model parameters, PDE regularity, and utility specifications. Full rigorous proofs would require meticulous tracking of these dependencies and bounds.
Our formal analysis (detailed in Appendix C) demonstrates that under suitable regularity conditions, particularly uniform parabolicity, a small Pontryagin first-order condition (FOC) gap ($\varepsilon$) achieved during the Stage 1 warm-up guarantees that the BPTT-derived costate estimates are $O(\varepsilon)$-close to the true optimal costates (up to the BPTT estimation error $\delta_{\mathrm{BPTT}}$). This provides a rigorous justification for the core idea of the P-PGDPO method: substituting these rapidly stabilizing and sufficiently accurate costate estimates into the analytical PMP formulas yields a near-optimal policy $\widehat{\pi}$ with an error bound of $O(\varepsilon+\delta_{\mathrm{BPTT}})$ (Theorem 1).
However, extending this rigorous proof framework to settings beyond uniform parabolicity, such as hypo-elliptic (H-class, (Hörmander, , 1967)) or boundary-degenerate (D-class, (Oleinik, , 2012)) models, or to general non-affine systems, faces significant analytical challenges, as key assumptions (like global $L^{p}$ regularity or the costate floor) may no longer hold.
Despite these formal hurdles, the fundamental algorithmic rationale underpinning the P-PGDPO approach remains persuasive for broader applicability. Neither Pontryagin’s Maximum Principle (PMP) (Pontryagin, , 2014; Yong and Zhou, , 2012) nor backpropagation-through-time (BPTT) (Goodfellow, , 2016) are inherently limited to uniformly parabolic or affine settings. The P-PGDPO strategy leverages the potentially advantageous decoupling of the problem into: (i) estimating potentially simpler or faster-converging costate functions via BPTT, and (ii) utilizing the robust analytical structure of PMP to construct the control.
This rationale, strongly supported by the numerical success demonstrated even in high dimensions for the affine benchmark case (Section 4), suggests that PG-DPO (especially the P-PGDPO variant) holds promise as a practical tool for a wider class of challenging continuous-time control problems. Nonetheless, confirming its effectiveness and potentially refining the theoretical guarantees for models outside the strictly regular class analyzed here remains an important avenue for future research, likely requiring tailored analysis or different analytical techniques (e.g., focusing on regions predominantly visited by the state process).
@@ 4 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance and scalability of the proposed PG-DPO framework (both Baseline and Projected PG-DPO (P-PGDPO) variants). We focus on the multi-asset, multi-factor Ornstein-Uhlenbeck (OU) portfolio optimization problem introduced in Section 2.3.2. The objective is to maximize the expected utility of terminal wealth, $J=\mathbb{E}[U(X_{T})]$, using a CRRA utility function with risk aversion $\gamma=2.0$. For these experiments, we specialize the general model by assuming no intermediate consumption (i.e., $C_{t}\equiv 0$) and focus solely on the utility from terminal wealth (effectively setting the bequest weight $\kappa=1$ if $U(X_{T})$ is the direct objective, or adjusting $\kappa$ appropriately if a specific HARA utility form with $a\neq 0$ was being precisely mapped).
This specific affine model setting allows for comparison against a semi-analytical benchmark solution derived from the HJB equation (details in Appendix B.2). Furthermore, to provide a comprehensive assessment against other contemporary deep learning techniques for high-dimensional control problems, we also compare our PG-DPO framework with the Deep BSDE method (Han et al., , 2018; Weinan, , 2017). The Deep BSDE approach reformulates the stochastic control problem in terms of a system of backward stochastic differential equations, approximating the unknown components of the solution (typically related to the value function gradient or control variables) using neural networks. These networks are trained by minimizing the error incurred in satisfying the BSDE dynamics and terminal conditions through Monte Carlo simulation.
We compare the policies generated by Baseline PG-DPO, Projected PG-DPO (P-PGDPO), and the Deep BSDE method against the HJB-based benchmark using the Root Mean Squared Error (RMSE) metric across various dimensions, scaling up to n=50 assets and k=10 state factors. For complete details on the specific parameter values, the implementations of PG-DPO and Deep BSDE (including network architectures and training setups), benchmark computation, and evaluation procedures, please refer to Appendix D. The subsequent subsections present and analyze the results of these comparative experiments.
@@ 4.1 Convergence Speed and Accuracy Analysis
Figure 1: Average policy RMSE comparison between Baseline PG-DPO (e.g., blue), Projected PG-DPO (P-PGDPO) (e.g., orange), and Deep BSDE (e.g., green) relative to the benchmark solution across different asset dimensions ($n$) with $k=10$ factors fixed. Values averaged over $k$ evaluation ‘slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation or visualization) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$. Further details on the evaluation methodology are provided in Appendix D. Note the logarithmic scale on the $y$-axis.
This section quantitatively compares the convergence speed and final accuracy of the proposed PG-DPO framework (Baseline and Projected PG-DPO (P-PGDPO) variants) against the Deep BSDE method. We analyze their performance across various dimensional settings, focusing on how the number of assets ($n$) and the number of state factors ($k$) impact accuracy. Figure 1 (Note: Assuming you will update the figure label in your main document if it includes Deep BSDE results) displays the average Root Mean Squared Error (RMSE) between the learned/derived policies and the benchmark optimal policy. The RMSE is plotted against training epochs on a logarithmic scale for configurations with $k=10$ state variables and varying numbers of assets ($n=1,10,50$).
The RMSE for P-PGDPO represents the error of the total portfolio policy $\bm{\pi}^{\mathrm{PMP}}$ (Eq. (11)), which combines myopic and hedging components. As detailed in Section 4.2, the myopic component’s error is typically negligible ($O(10^{-7})$). Thus, the total RMSE for P-PGDPO is almost entirely driven by the accuracy of estimating the hedging component using BPTT-derived costate derivatives, $\partial_{\mathbf{Y}}\lambda_{t}^{*}$.
Table 2 summarizes the minimum achieved policy RMSE and the epochs at which these minima were observed for the $k=10$ scenario with varying $n$. A consistent observation is the superior performance of the P-PGDPO method, which achieves significantly lower RMSE values compared to both Baseline PG-DPO and Deep BSDE across all tested asset dimensions. For instance, with $n=1,k=10$, P-PGDPO’s minimum RMSE is approximately an order of magnitude better than Deep BSDE and two orders of magnitude better than Baseline PG-DPO. Deep BSDE generally performs between the Baseline and P-PGDPO methods. While P-PGDPO maintains its lead as $n$ increases to 50, the performance gap between P-PGDPO and Deep BSDE tends to narrow, suggesting that the increasing dimensionality of $n$ poses challenges even for the PMP-based analytical solution, possibly due to the ill-conditioning of the covariance matrix $\Sigma$. All methods show an increase in RMSE as $n$ grows, indicating the inherent difficulty of high-dimensional asset allocation.
To further investigate the distinct effects of asset dimensionality ($n$) versus factor dimensionality ($k$), we conducted two additional sets of experiments.
Table 2: Summary of Minimum Policy RMSE and Convergence Epochs
@@
($n$ assets, $k=10$ factors)
n=1, k=10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
P-PGDPO $4.72\times 10^{-4}$ 1200
Deep BSDE $4.78\times 10^{-3}$ 600
n=10, k=10 Baseline PG-DPO $5.11\times 10^{-2}$ 9600
P-PGDPO $5.26\times 10^{-3}$ 1000
Deep BSDE $3.46\times 10^{-2}$ 1400
n=50, k=10 Baseline PG-DPO $1.05\times 10^{-1}$ 8800
P-PGDPO $1.58\times 10^{-2}$ 5400
Deep BSDE $4.90\times 10^{-2}$ 9200
@@
Note: Minimum Policy RMSE values are averaged across $k$ evaluation ’slices.’ For each slice $j=1,\dots,k$, the evaluation (e.g., for RMSE calculation) focuses on state factor $Y_{j}$, while other factors $Y_{l},l\neq j,$ are held constant at their respective long-term means $\theta_{Y,l}$. Further details on the evaluation methodology are provided in Appendix D. Epochs indicate the recorded iteration number when the minimum average RMSE was observed; evaluations and logging were performed every 200 iterations.
First, Table 3 presents results where the number of assets is fixed at $n=1$, and the number of state factors $k$ is varied ($k=1,5,10$). In this scenario, the performance of all three methodologies remains relatively stable across different values of $k$. For P-PGDPO, the RMSE even shows a slight improvement as $k$ increases, possibly due to more information being available for costate estimation. Deep BSDE and Baseline PG-DPO also exhibit no significant degradation in performance. This suggests that, for a small number of assets, the ”curse of dimensionality” associated with an increasing number of state factors ($k$) is largely managed by these neural network-based approaches. The underlying PMP structure or BSDE formulation appears robust to the increase in $k$ when $n$ is small and fixed, with Monte Carlo sampling and function approximation being the primary challenges rather than ill-conditioned linear algebra.
Table 3: Summary of Minimum Policy RMSE and Convergence Epochs for $n=1$ (Varying $k$)
$k$ factors Method Min. Policy RMSE Epochs at Min.
($n=1$ asset)
1 Baseline PG-DPO $3.26\times 10^{-2}$ 1000
P-PGDPO $1.20\times 10^{-3}$ 800
Deep BSDE $7.39\times 10^{-3}$ 800
5 Baseline PG-DPO $3.71\times 10^{-2}$ 1400
P-PGDPO $2.64\times 10^{-4}$ 800
Deep BSDE $1.02\times 10^{-2}$ 1400
10 Baseline PG-DPO $1.75\times 10^{-2}$ 2800
P-PGDPO $4.72\times 10^{-4}$ 1200
Deep BSDE $4.78\times 10^{-3}$ 600
Table 4: Summary of Minimum Policy RMSE and Convergence Epochs for $k=2$ (Varying $n$)
$n$ assets Method Min. Policy RMSE Epochs at Min.
($k=2$ factors)
1 Baseline PG-DPO $4.77\times 10^{-2}$ 1200
P-PGDPO $1.98\times 10^{-4}$ 800
Deep BSDE $4.63\times 10^{-3}$ 3400
10 Baseline PG-DPO $1.12\times 10^{-1}$ 4000
P-PGDPO $1.22\times 10^{-2}$ 1400
Deep BSDE $2.32\times 10^{-2}$ 2600
50 Baseline PG-DPO $2.24\times 10^{-1}$ 7200
P-PGDPO $5.55\times 10^{-2}$ 4400
Deep BSDE $1.88\times 10^{-1}$ 4400
Second, Table 4 shows results where the number of state factors is fixed at $k=2$, and the number of assets $n$ is varied ($n=1,10,50$). In stark contrast to the previous case, increasing $n$ leads to a significant deterioration in performance for all three methods. Baseline PG-DPO shows the most substantial increase in RMSE. While P-PGDPO still yields the lowest RMSE, its error also increases considerably with $n$. Deep BSDE’s performance also degrades, approaching that of Baseline PG-DPO for $n=50$. This pronounced degradation with increasing $n$ (even for a small $k$) points towards a different kind of challenge not solely related to the dimensionality of the state space for function approximation. A primary suspect is the numerical instability arising from the asset covariance matrix $\Sigma$. As $n$ increases, $\Sigma$ ($n\times n$) is more likely to become ill-conditioned (i.e., have a high condition number). The optimal policy calculation, particularly in the PMP framework (Eq. (11)), involves the inverse $\Sigma^{-1}$. If $\Sigma$ is ill-conditioned, any small errors in the estimated costates or market parameters can be greatly amplified by $\Sigma^{-1}$, leading to large errors in the computed policy. This suggests that the “curse” associated with large $n$ is not just about learning a higher-dimensional policy function but also about the increasing fragility of the underlying mathematical problem structure due to the properties of $\Sigma$. The problem itself becomes “damaged” or numerically challenging, impacting any method that relies on or implicitly reflects this structure.
These findings underscore that while neural network approaches have made significant strides in handling high-dimensional state spaces (related to $k$), the challenges posed by a large number of decision variables (assets $n$), especially when intertwined with ill-conditioned covariance matrices, require further attention. Strategies such as covariance matrix shrinkage or other regularization techniques applied prior to or within the optimization might be necessary to effectively address this “asset-dimension curse”.